In a shuffleboard game, the puck slides a total di
Answer #1

by newton's second law

force = mass * acceleration

frictional force = coefficient of friction * normal force

normal force = mg

frictional force = 0.2 * mg

soce force = ma

ma = 0.2 * mg

a = -0.2 * g

acceleration is negative since direction of force is opposite to the direction of motion

by third equation of motion

v^2 = u^2 + 2as

since the puck came to rest so v = 0 m/s

0 = u^2 - 2as

u = sqrt(2 * 0.2 * g * x)

u = sqrt(0.4 * g * x)

initial speed = sqrt(0.4 * g * x)